OpenStudy (anonymous):

A line passes through (4, –9) and (7, –2). a. Write an equation for the line in point-slope form. b. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

Equate the slope of each line that passes through the problem points. Solve the following for y:\[\frac{y+9}{x-4}=\frac{y+2}{x-7}\]\[y=\frac{7 x}{3}-\frac{55}{3} \]

OpenStudy (anonymous):

Standard Form:\[y=\frac{7 x}{3}-\frac{55}{3} \]\[y-\frac{7 x}{3}\text{=}-\frac{55}{3}\]\[3\left(y-\frac{7 x}{3}\right)\text{ = }\left(-\frac{55}{3}\right)3\]\[3 y-7 x\text{ =}-55\]